# Fixing Code Generation Errors for Large Language Models

### v1

@@ III-E2 Error Analysis Results
[nobreak=true] Answer to RQ3: Errors amenable to automated fixing should meet three crucial criteria: high consistency and predictability, high locality, and low complexity of fix. Based on these criteria, we identified three types of errors suitable for automated fixing including Missing Import, Function Overflow, and Inconsistent Indentation.
IV Methodology
RQ3 indicated that errors caused by three reasons can be directly and automatically fixed. This section details the proposed fixing method named LlmFix for such errors, which is further evaluated in Section V.
IV-A Overall Framework of LlmFix
Generally, LlmFix addresses errors in three distinct steps: Code Filtering (Step-1), Code Truncation (Step-2) and Importing Missing Modules (Step-3). Code Filtering focuses on rectifying Inconsistent Indentation, Code Truncation focuses on rectifying Function Overflow, while Importing Missing Modules concentrates on resolving Missing Import issues. Implementation details are described below. Generally, we have integrated the various fix methods mentioned previously to construct a complete fix process LlmFix, as shown in Fig. 9. To enhance understanding of this process, we have also created an example that demonstrates the specific function of each step within this fix process, as detailed in Fig. 10.
Figure 9: Flowchart of the method LlmFix.
Figure 10: An example of applying LlmFix. This example was crafted by us to help readers understand better. In practice, not all the fix steps are used within the same code segment.
IV-B Step-1: Code Filtering
In addition to the __main__ block, the code generated by the model may also contain input statements and assert statements. The former can cause the test program to get stuck waiting for user input, while the latter can lead to unnecessary errors since the test cases generated by the model itself might be incorrect. Therefore, we need to filter these out. Moreover, as shown in Listing LABEL:list_norm1 in Appendix, an unequal number of spaces used for indentation could cause an IndentationError. To avoid this issue, we replace all space-based indentations with tabs.
IV-C Step-2: Code Truncation
@@
IV-D Step-3: Importing Missing Modules
@@
V Evaluation
V-A Research Question
@@
RQ4: How effective is LlmFix? This RQ conducted comparative tests on the 14 LLMs tested in the empirical study, using both the HumanEval and MBPP datasets, under scenarios with and without the use of LlmFix. We also statistically analyzed the improvement LlmFix brings to different LLMs and the proportion of different error types it successfully fixes.
V-B Datasets
@@
We apply the LlmFix to run ten additional times on each of the 14 models selected in the empirical study using both the HumanEval and MBPP datasets’ test sets. This approach aims to demonstrate that the method is not only effective on HumanEval but also equally effective on the previously unstudied MBPP dataset, thus showcasing a certain degree of generalization capability. The MBPP (Mostly Basic Python Problems) dataset is a comprehensive collection of 974 Python programming tasks designed to evaluate code generation and program synthesis capabilities [60]. Each task in the dataset is accompanied by an average of 3 test cases, ensuring the functional correctness of the generated code. Prominent language models such as GPT-3 [26], Codex [13], and PaLM [61] have utilized MBPP for assessing their code generation proficiency.
As illustrated in Fig. 11, the MBPP dataset comprises six essential components: 1) task_id, a unique identifier assigned to each task, facilitating easy reference and organization of the code generation challenges. 2) text, a paragraph of text that provides generation requirements for guiding LLMs. 3) code, the standard solution for each task. 4) test_list, a testing list including several test cases. 5) test_setup_code, a piece of code that runs before executing the tests, ensuring that the tests are executed in the correct context. 6) challenge_test_list, a set of additional, more challenging test cases used to deeply evaluate the robustness and correctness of the generated code.
Figure 11: A schematic diagram of the first sample in MBPP test sets.
V-C Effectiveness of LlmFix in Error Fixing
@@
Table V lists the specific numerical characteristics of the ten tests conducted with LlmFix across 14 LLMs in the HumanEval dataset, including minimum, maximum, mean, and standard deviation. Comparing with Table I, we found that with LlmFix, the average test results of each model improved by about 9.5%, with a maximum improvement of 16.46% percentage points. In terms of a total number of errors, LlmFix addressed 17.07% of all errors in HumanEval (reducing errors from 12,837 to 10,646). Additionally, out of the ten models that publicly reported their performance, six of them surpassed the reported SOTA performance using our fixing process. We performed a Wilcoxon signed-rank test on the mean test results before and after fixing, obtaining a p-value of $1.22\times 10^{-4}$, which is much smaller than the significance level of 0.05. This result indicates that our fixing process significantly improved the performance of the models. Furthermore, the standard deviation (STD) of all models is much lower than their average value (less than 10% of the average), indicating very low data variability.
@@
TABLE V: Accuracy of ten tests (with LlmFix) of LLMs on HumanEval dataset in terms of minimum, maximum, mean, and standard deviation (STD) of Pass@1 values.
@@
We have identified that among all 19 error causes, only three causes are amenable to automatic fixing: Missing Imports, Function Overflow, and Inconsistent Indentation. Among them, Missing Import is one of the causes leading to NameError, Function Overflow is one of the causes leading to SyntaxError, and Inconsistent Indentation is one of the causes leading to IndentationError. Given that compiling statistics based on error types is more convenient, and considering the low frequency of IndentationErrors and their similarity to SyntaxErrors in terms of root causes [62], we have incorporated the IndentationError statistics into the SyntaxError category here. Subsequently, we compared the frequency of NameErrors and SyntaxErrors before and after applying LlmFix. The results of this comparison are illustrated in Fig. 12. We found that our proposed fixing process fixed an average of 90.26% of NameErrors and 99.74% of SyntaxErrors.
@@
Figure 12: The average sample number of all NameErrors and SyntaxErrors (including IndentationErrors) per round of tests using the HumanEval dataset, without or with LlmFix.
@@
Table VI shows the specific numerical characteristics of the ten tests conducted across 14 LLMs in the MBPP dataset, comparing scenarios with and without LlmFix, including minimum, maximum, mean, and standard deviation. We found that with LlmFix, the average test results of each model improved by about 5.4%, with a maximum improvement of 8.48%, on average. We performed a Wilcoxon signed-rank test on the mean test results before and after fixing, obtaining a p-value of $1.22\times 10^{-4}$, which is much smaller than the significance level of 0.05. This result indicates that LlmFix significantly improved the performance of the models. Furthermore, the standard deviation (STD) of all models is much lower than their average value (less than 10% of the average), indicating very low data variability.
Model (Without LlmFix) Min Max Mean STD
CodeLlama-Python-7B 40.60 43.40 42.12 0.94
CodeLlama-Python-13B 45.60 48.40 46.86 0.92
CodeLlama-Python-34B 47.80 51.00 49.82 1.03
WizardCoder-Python-V1.0-7B 48.60 51.40 50.08 0.92
WizardCoder-Python-V1.0-13B 49.80 52.20 50.76 0.79
WizardCoder-Python-V1.0-34B 55.60 57.40 56.68 0.53
Phind-CodeLlama-V2-34B 56.20 58.80 57.16 0.91
Phi-1-1.3B 46.20 47.80 47.02 0.55
Phi-1.5-1.3B 34.00 38.60 36.40 1.46
Phi-2-2.7B 41.80 44.00 42.70 0.73
GPT-3.5-turbo 63.20 66.40 64.64 1.02
CodeTulu-2-7B 43.60 45.40 44.50 0.63
CodeTulu-2-13B 47.40 50.20 49.02 0.80
CodeTulu-2-34B 53.00 56.40 54.38 0.92
Model (With LlmFix) Min Max Mean STD
CodeLlama-Python-7B 47.20 49.80 48.04 0.82
CodeLlama-Python-13B 51.00 54.20 52.46 1.01
CodeLlama-Python-34B 56.20 58.20 57.24 0.74
WizardCoder-Python-V1.0-7B 54.00 55.40 54.42 0.47
WizardCoder-Python-V1.0-13B 53.00 54.80 54.00 0.63
WizardCoder-Python-V1.0-34B 60.00 62.60 61.12 0.83
Phind-CodeLlama-V2-34B 60.20 63.20 61.44 0.85
Phi-1-1.3B 47.60 50.80 49.32 0.95
Phi-1.5-1.3B 42.40 45.00 43.86 0.85
Phi-2-2.7B 50.00 53.00 51.18 0.91
GPT-3.5-turbo 70.40 73.40 72.18 0.87
CodeTulu-2-7B 46.80 51.20 49.36 1.43
CodeTulu-2-13B 51.40 55.40 53.52 1.21
CodeTulu-2-34B 57.80 60.60 59.30 0.86
@@
TABLE VI: Accuracy of ten tests(without/with LlmFix) of LLMs on MBPP dataset in terms of minimum, maximum, mean, and standard deviation (STD) of Pass@1 values.
@@
Figure 13: The average sample number of all NameErrors and SyntaxErrors (including IndentationErrors) per round of tests using the MBPP dataset, without or with LlmFix.
@@
[nobreak=true] Answer to RQ4: Our proposed method, LlmFix, demonstrates the capability to resolve nearly all errors that are amenable to automated fixing in both the HumanEval and MBPP datasets. LlmFix can provide 9.5% and 5.4% average accuracy improvements for 14 LLMs on HumanEval and MBPP, respectively.
VI Discussion
VI-A Implication
Based on our experimental results and discussion, this section provides implications for model developers and future research.
For model developers, when introducing a new LLM, it is beneficial to concurrently develop a customized version of LlmFix, thereby further enhancing the model’s practical performance. We developed a general-purpose LlmFix based on the experimental results of 14 LLMs on the HumanEval dataset, as discussed in Section V. This version has demonstrated excellent performance on both the HumanEval and MBPP datasets, thus serving as a benchmark tool applicable to various models and datasets. However, to achieve better results, model developers can fine-tune the baseline LlmFix specifically for their models, potentially leading to more significant performance improvements for those particular models.
For future research directions, LlmFix can serve as a novel metric for assessing code quality. We can posit that code passing tests without requiring LlmFix processing is of higher quality than code needing LlmFix treatment to pass tests. As mentioned in subsection IV-A, LlmFix comprises three steps, though not all code requires all three steps as illustrated in Fig. 10. Building on this observation, we can further infer that when $m<n$, code passing tests after only $m$ steps of LlmFix processing is of higher quality than code requiring $n$ steps to pass tests. This provides a new perspective for a more nuanced evaluation of code quality.
VI-B Threats to Validity
There are some potential threats affecting the validity of our experimental results and conclusions.
Limited Dataset. In our research, we evaluated the code generation accuracy of 14 selected models. Due to our limited computation resources, we focused our testing on two widely-used datasets: HumanEval and MBPP. While these datasets are well-established in the field, the narrow scope of our data selection raises questions about the generalizability of our experimental results and findings. The limited dataset diversity leaves room for uncertainty regarding how our conclusions might apply to other, potentially more diverse, or specialized, coding datasets. This limitation underscores the need for future studies to incorporate a broader range of datasets to validate and potentially extend our findings.
Limited Testing. Our study focused on testing the Pass@1 results of 14 LLMs on the HumanEval and MBPP datasets, leading to the development and validation of LlmFix. However, due to constraints in cost and time, we were unable to conduct tests for Pass@10 or even Pass@100. Extending our experiments to include these higher sampling numbers would undoubtedly make our study more comprehensive, thereby strengthening the persuasiveness and reliability of our conclusions. This highlights a potential area for improvement in future research.
VII Conclusion

### v2

@@ 3.5.2. Error Analysis Results
[nobreak=true] Answer to RQ3: We identified three types of errors suitable for automated fixing including Missing Import, Function Overflow, and Inconsistent Indentation because these errors possess high consistency and predictability, high locality, and low fixing complexity.
4. Methodology
RQ3 indicated that errors caused by three reasons can be directly and automatically fixed. This section details the proposed fixing method named LlmFix for such errors, which is further evaluated in Section 5.
4.1. Overall Framework of LlmFix
Generally, LlmFix addresses errors in three distinct steps: Code Filtering (Step-1), Code Truncation (Step-2) and Importing Missing Modules (Step-3). Code Filtering focuses on rectifying Inconsistent Indentation, Code Truncation focuses on rectifying Function Overflow, while Importing Missing Modules concentrates on resolving Missing Import issues. Implementation details are described below. Generally, we have integrated the various fix methods mentioned previously to construct a complete fix process LlmFix, as shown in Fig. 5.
Figure 5. Flowchart of the method LlmFix.
4.2. Step-1: Code Filtering
In addition to the __main__ block, the code generated by the model may also contain input statements and assert statements. The former can cause the test program to get stuck waiting for user input, while the latter can lead to unnecessary errors since the test cases generated by the model itself might be incorrect. Therefore, we need to filter these out. Moreover, as shown in Listing LABEL:list_norm1 in Appendix, an unequal number of spaces used for indentation could cause an IndentationError. To address this issue, we first standardized all indentations by replacing them with tab characters. Subsequently, we further adjusted the indentations based on the code’s logical structure. The detailed procedure is illustrated in Algorithm 1.
Algorithm 1 Pseudocode for Resolving Inconsistent Indentation
1:Source code $SourceCode$
2:Correctly indented code compared to $SourceCode$
3:Initialization: Read the code and store each line in a list
4:$lines\leftarrow$ SplitLines($SourceCode$)
5:First Pass - Normalize indentation sizes
6:for $i\leftarrow 0$ to length($lines$) $-1$ do
7:    $line\leftarrow lines[i]$
8:    $indentSize\leftarrow$ CountLeadingSpaces($line$)
9:    if $indentSize\mod 4=0$ then
10:       Replace leading spaces with a tab in $lines[i]$
11:    else
12:       if $indentSize\mod 2=0$ then
13:          $newIndent\leftarrow indentSize-2$
14:          Replace leading spaces with $newIndent$ spaces in $lines[i]$
15:          Replace leading spaces with a tab in $lines[i]$
16:       else
17:          if $indentSize\mod 4<2$ then
18:             $newIndent\leftarrow indentSize-(indentSize\mod 4)$
19:          else
20:             $newIndent\leftarrow indentSize+(4-(indentSize\mod 4))$
21:          end if
22:          Replace leading spaces with $newIndent$ spaces in $lines[i]$
23:          Replace leading spaces with a tab in $lines[i]$
24:       end if
25:    end if
26:end for
27:Second Pass - Ensure logical code structure based on colons
28:for $i\leftarrow 0$ to length($lines$) $-2$ do
29:    $currentLine\leftarrow lines[i]$
30:    $nextLine\leftarrow lines[i+1]$
31:    if EndsWithColon($currentLine$then
32:       $currentIndent\leftarrow$ CountLeadingTabs($currentLine$)
33:       $nextIndent\leftarrow$ CountLeadingTabs($nextLine$)
34:       if $nextIndent=currentIndent$ then
35:          Increase indentation of $nextLine$ by one tab
36:          Update $lines[i+1]$ with the new indentation
37:       end if
38:    end if
39:end for
4.3. Step-2: Code Truncation
@@
4.4. Step-3: Importing Missing Modules
@@
5. Evaluation
5.1. Research Question
The evaluation aims to assess the superiority of LlmFix over other error-fixing methods in code error-fixing tasks, as well as whether applying it to function-level code generation tasks can achieve significant improvements compared to not using the method.
RQ4: How effective is LlmFix for error fixing? This RQ constructs a dataset, LlmErrorEval, based on the 12k+ errors collected in RQ2 to evaluate the performance of existing fixing methods. Using this dataset, we assess the fix rate of five existing static analysis tools, such as CARDUMEN (Martinez and Monperrus, 2018) and KALI (Qi et al., 2015), which are parts of the recent fixing framework FixKit (Smytzek et al., 2024), as well as our static analysis tool LlmFix and two recent fixing LLMs: LDB (Zhong et al., 2024) and CoR (Wang et al., 2023). Additionally, we integrate LlmFix with LDB and CoR to evaluate their combined performance in terms of fix rate and time consumption.
@@ 5.1. Research Question
RQ5: How effective is LlmFix for improving function-level code generation of foundation LLMs? This RQ conducts comparative tests on the 14 LLMs tested in the empirical study, using both the HumanEval and MBPP datasets, under scenarios with and without the use of LlmFix. We also statistically analyze the improvement LlmFix brings to different LLMs and the proportion of different error types it successfully fixes.
5.2. RQ4: How Effective is LlmFix for Error Fixing?
5.2.1. Existing Fixing Methods
We compare LlmFix with seven existing fixing methods:
  • GENPROG (Weimer et al., 2009), a static analysis method that evolves program variants based on genetic programming.
  • MUTREPAIR (Debroy and Wong, 2010), a static analysis method that generates patches through mutation operators.
  • KALI (Qi et al., 2015), a static analysis method that removes or modifies faulty code to create patches.
  • CARDUMEN (Martinez and Monperrus, 2018), a static analysis method that uses probabilistic models to generate fix templates.
  • AE (Weimer et al., 2013), a static analysis method that leverages program equivalence relations for patch generation.
  • CoR (Wang et al., 2023), an LLM-based fixing method that iteratively interprets errors and applies fixes using natural language.
  • LDB (Zhong et al., 2024), an LLM-based fixing method that incorporates runtime execution information to improve code fixes.
5.2.2. LlmErrorEval
Based on the 12k+ errors collected in RQ2, we build a dataset called LlmErrorEval for evaluating fixing performance. The dataset consists of seven parts: 1) prompt, a function specification provided as input for LLMs; 2) erroneous_code, a code snippet containing errors generated by LLMs; 3) LLM, the name of the model that generated the erroneous code, such as CodeLlama-Python-7B; 4) error_type, the type of error, such as AssertionError; 5) error_message, a brief description of the error, output by the compiler; 6) test_cases, a set of test cases used to evaluate the functionality of the code; 7) canonical_solution, the standerd solution for the given prompt.
5.2.3. Comparative Experimental Results on LlmErrorEval
Using LlmErrorEval, we evaluated the fix rates of the seven methods mentioned above and further tested the performance when combining the LLM-based fixing methods, LDB (Zhong et al., 2024) and CoR (Wang et al., 2023), with LlmFix. Table 4 presents the number of errors successfully fixed by each method and their corresponding fix rates. The results reveal that current static analysis tools perform poorly, fixing only an average of 92 errors (0.72% of the total 12,837), which aligns with the findings of Fan et al. (Fan et al., 2023). In contrast, LLM-based fixing methods show significantly better performance, fixing an average of 1,780 errors (13.87%). LlmFix, as a static analysis method, fixed 2,191 errors (17.07%), demonstrating its superiority among similar methods. Furthermore, it complements existing LLM-based fixing methods: combining LlmFix with CoR or LDB achieves even better results, fixing an average of 2,991 errors (23.30%). We also measured the time required for LlmFix to fix errors. The results show that LlmFix takes an average of only 11.50ms per fix, indicating that combining it with other methods does not introduce additional computational overhead.
Table 4. Evaluation results on LlmErrorEval.
Methods Errors Fixed Fix Rate Methods Errors Fixed Fix Rate
GENPROG 93 0.72% CoR 1547 12.05%
MUTREPAIR 86 0.67% LDB 2012 15.67%
KALI 95 0.74% LlmFix 2191 17.07%
CARDUMEN 94 0.73% LlmFix+CoR 2784 21.69%
AE 93 0.72% LlmFix+LDB 3198 24.91%
{mdframed}
[nobreak=true] Answer to RQ4: We constructed an evaluation dataset, LlmErrorEval, based on the 12k+ errors collected in RQ2. Experimental results show that existing static analysis methods exhibit limited effectiveness, achieving only a 0.72% fix rate. In contrast, LlmFix achieves a significantly higher fix rate of 17.1%, outperforming the state-of-the-art LDB by 8.9%. Furthermore, integrating LlmFix with LDB raises the fix rate to 24.9%, with minimal additional computational overhead, as LlmFix requires an average of only 11.5ms per fix.
5.3. RQ5: How Effective is LlmFix for Improving Function-Level Code Generation of Foundation LLMs?
5.3.1. Dataset
@@
We apply the LlmFix to run ten additional times on each of the 14 models selected in the empirical study using both the HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) datasets’ test sets. This approach aims to demonstrate that the method is not only effective on HumanEval but also equally effective on the previously unstudied MBPP dataset, thus showcasing a certain degree of generalization capability. The MBPP (Mostly Basic Python Problems) dataset is a comprehensive collection of 974 Python programming tasks designed to evaluate code generation and program synthesis capabilities (Google-Search, 2021). Each task in the dataset is accompanied by an average of 3 test cases, ensuring the functional correctness of the generated code. Prominent language models such as GPT-3 (Brown et al., 2020), Codex (Chen et al., 2021), and PaLM (Chowdhery et al., 2023) have utilized MBPP for assessing their code generation proficiency. The MBPP dataset comprises the same essential components as HumanEval as described in Section 3.2.
5.3.2. Effectiveness of LlmFix in Error Fixing
@@
Table 5 lists the specific numerical characteristics of the ten tests conducted with LlmFix across 14 LLMs in the HumanEval dataset, including minimum, maximum, mean, and standard deviation. Comparing with Table 1, we found that with LlmFix, the average test results of each model improved by about 9.5%, with a maximum improvement of 16.46% percentage points. Meanwhile, it is worth noting that LlmFix does not have any negative impact on any LLM, as it only attempts to fix code that contains errors. Additionally, out of the ten models that publicly reported their performance, six of them surpassed the reported SOTA performance using our fixing process. We performed a Wilcoxon signed-rank test on the mean test results before and after fixing, obtaining a p-value below 0.05. This result indicates that our fixing process significantly improved the performance of the models. Furthermore, the standard deviation (STD) of all models is much lower than their average value (less than 10% of the average), indicating very low data variability.
@@
Table 5. Accuracy of ten tests (with LlmFix) of LLMs on HumanEval dataset in terms of minimum, maximum, mean, and standard deviation (STD) of Pass@1 values.
@@
We have identified that among all 19 error causes, only three causes are amenable to automatic fixing: Missing Imports, Function Overflow, and Inconsistent Indentation. Among them, Missing Import is one of the causes leading to NameError, Function Overflow is one of the causes leading to SyntaxError, and Inconsistent Indentation is one of the causes leading to IndentationError. Given that compiling statistics based on error types is more convenient, and considering the low frequency of IndentationErrors and their similarity to SyntaxErrors in terms of root causes (Python, 2024), we have incorporated the IndentationError statistics into the SyntaxError category here. Subsequently, we compared the frequency of NameErrors and SyntaxErrors before and after applying LlmFix. The results of this comparison are illustrated in Fig. 6. We found that our proposed fixing process fixed an average of 90.26% of NameErrors and 99.74% of SyntaxErrors.
@@
Figure 6. The average sample number of all NameErrors and SyntaxErrors (including IndentationErrors) per round of tests using the HumanEval dataset, without or with LlmFix.
@@
Table LABEL:tab_validation shows the specific numerical characteristics of the ten tests conducted across 14 LLMs in the MBPP dataset, comparing scenarios with and without LlmFix, including minimum, maximum, mean, and standard deviation. We found that with LlmFix, the average test results of each model improved by about 5.4%, with a maximum improvement of 8.48%, on average. We performed a Wilcoxon signed-rank test on the mean test results before and after fixing, obtaining a p-value $<$ 0.05. This result indicates that LlmFix significantly improved the performance of the models. Furthermore, the standard deviation (STD) of all models is much lower than their average value (less than 10% of the average), indicating very low data variability.
@@
Table 6. Accuracy of ten tests(without/with LlmFix) of LLMs on MBPP dataset in terms of minimum, maximum, mean, and standard deviation (STD) of Pass@1 values.
Model Performance without LlmFix Performance with LlmFix
Min Max Mean STD Min Max Mean STD
CodeLlama-Python-7B 40.60 43.40 42.12 0.94 47.20 49.80 48.04 0.82
CodeLlama-Python-13B 45.60 48.40 46.86 0.92 51.00 54.20 52.46 1.01
CodeLlama-Python-34B 47.80 51.00 49.82 1.03 56.20 58.20 57.24 0.74
WizardCoder-Python-V1.0-7B 48.60 51.40 50.08 0.92 54.00 55.40 54.42 0.47
WizardCoder-Python-V1.0-13B 49.80 52.20 50.76 0.79 53.00 54.80 54.00 0.63
WizardCoder-Python-V1.0-34B 55.60 57.40 56.68 0.53 60.00 62.60 61.12 0.83
Phind-CodeLlama-V2-34B 56.20 58.80 57.16 0.91 60.20 63.20 61.44 0.85
Phi-1-1.3B 46.20 47.80 47.02 0.55 47.60 50.80 49.32 0.95
Phi-1.5-1.3B 34.00 38.60 36.40 1.46 42.40 45.00 43.86 0.85
Phi-2-2.7B 41.80 44.00 42.70 0.73 50.00 53.00 51.18 0.91
GPT-3.5-turbo 63.20 66.40 64.64 1.02 70.40 73.40 72.18 0.87
CodeTulu-2-7B 43.60 45.40 44.50 0.63 46.80 51.20 49.36 1.43
CodeTulu-2-13B 47.40 50.20 49.02 0.80 51.40 55.40 53.52 1.21
CodeTulu-2-34B 53.00 56.40 54.38 0.92 57.80 60.60 59.30 0.86
@@
Figure 7. The average sample number of all NameErrors and SyntaxErrors (including IndentationErrors) per round of tests using the MBPP dataset, without or with LlmFix.
@@
[nobreak=true] Answer to RQ5: Evaluations of LlmFix on both the HumanEval and MBPP datasets demonstrated its effectiveness, successfully resolving all directly fixable errors and improving the Pass@1 performance of 14 LLMs by an average of 7.5%.
6. Discussion
6.1. Ablation Study
Table 7. Results of our ablation study. CF, CT, and IMM represent the three steps in LlmFix: Code Filtering, Code Truncation, and Importing Missing Modules, respectively.
Model (on HumanEval) LlmFix w/o CF LlmFix w/o CT LlmFix w/o IMM LlmFix
CodeLlama-Python-7B 33.60 ($\downarrow$15.75%) 37.62 ($\downarrow$05.67%) 35.92 ($\downarrow$09.93%) 39.88
CodeLlama-Python-13B 38.11 ($\downarrow$14.38%) 41.58 ($\downarrow$06.58%) 40.30 ($\downarrow$09.46%) 44.51
CodeLlama-Python-34B 43.96 ($\downarrow$19.53%) 47.93 ($\downarrow$12.26%) 47.68 ($\downarrow$12.72%) 54.63
WizardCoder-Python-V1.0-7B 43.99 ($\downarrow$14.62%) 47.74 ($\downarrow$07.34%) 46.09 ($\downarrow$10.54%) 51.52
WizardCoder-Python-V1.0-13B 53.64 ($\downarrow$04.06%) 54.49 ($\downarrow$02.54%) 54.27 ($\downarrow$02.93%) 55.91
WizardCoder-Python-V1.0-34B 60.45 ($\downarrow$10.36%) 66.19 ($\downarrow$01.85%) 60.03 ($\downarrow$10.99%) 67.44
Phind-CodeLlama-V2-34B 65.30 ($\downarrow$08.54%) 68.47 ($\downarrow$04.10%) 66.22 ($\downarrow$07.25%) 71.40
Phi-1-1.3B 46.04 ($\downarrow$09.80%) 49.70 ($\downarrow$02.63%) 48.41 ($\downarrow$05.15%) 51.04
Phi-1.5-1.3B 29.70 ($\downarrow$26.54%) 34.88 ($\downarrow$13.73%) 31.10 ($\downarrow$23.08%) 40.43
Phi-2-2.7B 40.36 ($\downarrow$22.40%) 42.31 ($\downarrow$18.65%) 43.96 ($\downarrow$15.48%) 52.01
GPT-3.5-turbo 68.23 ($\downarrow$06.05%) 72.56 ($\downarrow$00.08%) 62.80 ($\downarrow$13.52%) 72.62
CodeTulu-2-7B 36.04 ($\downarrow$17.79%) 40.55 ($\downarrow$07.50%) 37.68 ($\downarrow$14.05%) 43.84
CodeTulu-2-13B 41.70 ($\downarrow$17.51%) 45.79 ($\downarrow$09.42%) 42.56 ($\downarrow$15.81%) 50.55
CodeTulu-2-34B 49.33 ($\downarrow$10.41%) 53.84 ($\downarrow$02.22%) 51.46 ($\downarrow$06.54%) 55.06
Model (on MBPP) LlmFix w/o CF LlmFix w/o CT LlmFix w/o IMM LlmFix
CodeLlama-Python-7B 42.30 ($\downarrow$11.95%) 42.36 ($\downarrow$11.82%) 43.66 ($\downarrow$09.12%) 48.04
CodeLlama-Python-13B 47.36 ($\downarrow$09.72%) 46.92 ($\downarrow$10.56%) 46.58 ($\downarrow$11.21%) 52.46
CodeLlama-Python-34B 50.80 ($\downarrow$11.25%) 49.80 ($\downarrow$13.00%) 51.74 ($\downarrow$09.61%) 57.24
WizardCoder-Python-V1.0-7B 51.10 ($\downarrow$06.10%) 50.32 ($\downarrow$07.53%) 53.18 ($\downarrow$02.28%) 54.42
WizardCoder-Python-V1.0-13B 51.12 ($\downarrow$05.33%) 51.12 ($\downarrow$05.33%) 52.88 ($\downarrow$02.07%) 54.00
WizardCoder-Python-V1.0-34B 56.68 ($\downarrow$07.26%) 56.60 ($\downarrow$07.40%) 58.84 ($\downarrow$03.73%) 61.12
Phind-CodeLlama-V2-34B 57.44 ($\downarrow$06.51%) 57.28 ($\downarrow$06.77%) 58.34 ($\downarrow$05.05%) 61.44
Phi-1-1.3B 47.02 ($\downarrow$04.66%) 47.02 ($\downarrow$04.66%) 49.22 ($\downarrow$00.20%) 49.32
Phi-1.5-1.3B 39.94 ($\downarrow$08.94%) 38.50 ($\downarrow$12.22%) 39.76 ($\downarrow$09.35%) 43.86
Phi-2-2.7B 45.26 ($\downarrow$11.57%) 44.06 ($\downarrow$13.91%) 46.92 ($\downarrow$08.32%) 51.18
GPT-3.5-turbo 64.64 ($\downarrow$10.45%) 64.40 ($\downarrow$10.78%) 71.00 ($\downarrow$01.63%) 72.18
CodeTulu-2-7B 45.42 ($\downarrow$07.98%) 48.04 ($\downarrow$02.67%) 45.10 ($\downarrow$08.63%) 49.36
CodeTulu-2-13B 49.02 ($\downarrow$08.41%) 51.60 ($\downarrow$03.59%) 49.08 ($\downarrow$08.30%) 53.52
CodeTulu-2-34B 54.38 ($\downarrow$08.30%) 54.38 ($\downarrow$08.30%) 54.58 ($\downarrow$07.96%) 59.30
Table 7 presents the results of our ablation study. Statistically, the mean decreases are 11.3%, 7.6%, and 8.7%, with standard deviations of 0.055, 0.045, and 0.051, when LlmFix is used without Code Filtering, Code Truncation, or Importing Missing Modules, respectively. These statistics suggest that the three components of LlmFix have varying levels of impact, and the high variability is evident as the ratios of standard deviations to mean values exceed 10%. For example, on the HumanEval dataset, the performance decrease is 26.54% on Phi-1.5-1.3B but only 4.06% on WizardCoder-Python-V1.0-13B when Code Filtering is removed. This highlights that different LLMs have significantly different sensitivities to each component of LlmFix.
6.2. Performance on Recent Foundation LLMs
To further investigate the effectiveness of LlmFix, we introduced six recent open-source foundation LLMs derived from three families: 1) DeepSeek-Coder-Instruct (1.3B/6.7B/33B) (Guo et al., 2024); 2) Llama-3.1-Instruct (8B) (AI, 2024a), proposed by Meta AI; 3) Llama-3.2-Instruct (1B/3B) (AI, 2024b), proposed by Meta AI; Experimental results show an average performance increase of 12.4% with LlmFix, where Llama-3.2-1B-Instruct achieved the highest improvement of 22.53% on MBPP, while DeepSeek-Coder-1.3B-Instruct on MBPP showed the lowest improvement of 6.92%.
Table 8. The test results of six recent foundation LLMs with and without LlmFix.
Model Baseline LlmFix
HumanEval MBPP HumanEval MBPP
DeepSeek-Coder-1.3B-Instruct 53.23 48.82 60.37 ($\uparrow$13.41%) 52.20 ($\uparrow$06.92%)
DeepSeek-Coder-6.7B-Instruct 62.44 58.78 70.12 ($\uparrow$12.30%) 67.60 ($\uparrow$15.01%)
DeepSeek-Coder-33B-Instruct 59.52 62.92 69.51 ($\uparrow$16.78%) 68.80 ($\uparrow$09.35%)
Llama-3.1-8B-Instruct 57.99 56.22 64.63 ($\uparrow$11.45%) 61.00 ($\uparrow$08.50%)
Llama-3.2-1B-Instruct 32.14 30.36 35.37 ($\uparrow$10.05%) 37.20 ($\uparrow$22.53%)
Llama-3.2-3B-Instruct 46.46 47.66 51.83 ($\uparrow$11.56%) 52.80 ($\uparrow$10.78%)
6.3. Implication
For future research directions, the LlmErrorEval, as described in Section 5.2, could serve as a benchmark. Besides, the performance of LlmFix combined with LDB could provide a strong baseline for evaluating new fixing methods. Developing new fixing methods, particularly static analysis methods, will be essential for future LLMs due to their efficiency and effectiveness. For instance, new LLM developers could fine-tune a specific version of LlmFix tailored to their LLMs’ characteristics and evaluate its performance in combination with other LLM-based fixing methods, such as LDB, on the LlmErrorEval benchmark. This approach would allow developers to assess the effectiveness of their methods and ultimately enhance the code generation capabilities of their LLMs.
6.4. Limitation
This study only implemented the python version of LlmFix, which limits its practicality. For other programming languages, the predefined rules in LlmFix may no longer apply due to differences in syntax. However, the underlying principles remain consistent, and the rules can be adapted to suit different programming languages. For example, when generating code in languages like C++ or Java, issues such as Missing Import or Function Overflow may still occur. By leveraging the principles of LlmFix, the predefined rules can be modified accordingly. Specifically, in the Importing Missing Modules component that addresses the Missing Import issue, the Python-specific rules that add statements like import XXX or from XXX import XXX can be replaced with the appropriate C++ syntax, such as adding #include <XXX> statements.
7. Conclusion